# PaDeLLM-NER: Parallel Decoding in Large Language Models for Named Entity Recognition

### v4

@@ Abstract
In this study, we aim to reduce generation latency for Named Entity Recognition (NER) with Large Language Models (LLMs). The main cause of high latency in LLMs is the sequential decoding process, which autoregressively generates all labels and mentions for NER, significantly increase the sequence length. To this end, we introduce Parallel Decoding in LLM for NER (PaDeLLM-NER), a approach that integrates seamlessly into existing generative model frameworks without necessitating additional modules or architectural modifications. PaDeLLM-NER allows for the simultaneous decoding of all mentions, thereby reducing generation latency. Experiments reveal that PaDeLLM-NER significantly increases inference speed that is 1.76 to 10.22 times faster than the autoregressive approach for both English and Chinese. Simultaneously it maintains the quality of predictions as evidenced by the performance that is on par with the state-of-the-art across various datasets.
[…]
PaDeLLM-NER: Parallel Decoding in Large Language Models for Named Entity Recognition
**footnotetext: Corresponding author.
@@ 1 Introduction
Augmented Language Paolini et al. (2020); Das et al. (2023)
Japan, co-hosts of the World Cup in 2002 and ranked 20th in the world by FIFA, are favourites to regain their title here.
[Japan — LOC], co-hosts of the [World Cup — MISC] in 2002 and ranked 20th in the world by [FIFA — ORG], are favourites to regain their title here.
Structured Annotation Lu et al. (2022, 2023a); Wang et al. (2023c)
Cuttitta announced his retirement after the 1995 World Cup , where he took issue with being dropped from the Italy side that faced England in the pool stages.
((PER): (Cuttitta), (MISC): (1995 World Cup), (LOC): (Italy), (LOC): (England), (ORG): (NULL))
@@
Named Entity Recognition (NER), a fundamental task in Natural Language Processing (NLP), aims to extract structured information from unstructured text data. This includes identifying and categorizing key elements such as Organization, Geopolitical Entity and so on (referred to as “labels”) in inputs, and pairing them with relevant text spans extracted from the text (termed “mentions”). Conventionally, NER tasks are carried out through an extractive paradigm that entails token-level classification and the subsequent extraction of identified tokens Ma et al. (2020); Liu et al. (2021).
@@
Recent advancements in Large Language Models (LLMs) Raffel et al. (2020a); Muennighoff et al. (2022); Touvron et al. (2023a, b); Bai et al. (2023); Gao et al. (2023); Liu et al. (2023d); Yang et al. (2023a) have revolutionized numerous foundational tasks in NLP, including NER tasks Paolini et al. (2020); Lu et al. (2022); Das et al. (2023); Lu et al. (2023a); Wang et al. (2023c), through the adoption of a generative paradigm. This paradigm involves instruction-tuning a sequence-to-sequence (seq2seq) model. The model takes a sequence of unstructured text as input and produces a sequence of structured label-mention pairs as output. Generally, the output structured string should be formatted to meet two criteria: (1) it should have a clear and straightforward structure that facilitates post-processing for label and mention extraction, and (2) it needs to be generated fluidly and efficiently from the perspective of language models Wang et al. (2023b).
@@
In Table 1, we list two typically used autoregressive output formats found in the literature : (1) accommodate original input text to contain label information, which is referred to as “augmented language” Paolini et al. (2020); Das et al. (2023); (2) directly using a customized, easily-parsed structured format to output all labels and mentions, which is called “structured annotation” Lu et al. (2022, 2023a); Wang et al. (2023c). These formats present certain challenges. For example, augmented language necessitates duplicating all original input text, thereby increasing output length and resulting in inference inefficiency. While structure annotation avoids replicating the entire input, it produces all labels and mentions in an autoregressive manner. This implies that each subsequently generated pair depends on its preceding pairs, and when the number of label-mention pairs is large, it will lead to longer sequences. As demonstrated in Chen et al. (2023c); Ning et al. (2023), high latency in LLMs mainly stems from lengthy sequence generation, we believe that by reducing the length of sequence, a more efficient inference scheme can be provided for NER tasks.
@@
In light of this, we propose Parallel Decoding in LLM for NER (PaDeLLM-NER), a novel approach to accelerate the inference of NER tasks for LLMs. PaDeLLM-NER empowers the model with the capability to predict a single label-mention pair within a single sequence, subsequently aggregating all sequences to generate the final NER outcome. Specifically, in the training phase, we reconstruct the instruction tuning tasks, enabling LLMs to predict the count of mentions for a specific label and to identify the $n^{th}$ mention within the entire input for that label (Figure 1). In the inference phase, LLMs first predict the number of mentions for all labels, then predict all label-mention pairs in parallel (Figure 2). Finally, results from all sequences are aggregated and duplicate mentions across labels are eliminated based on prediction probability. This approach results in a more efficient inference method for NER tasks, producing shorter output strings and enabling parallel decoding label-mention pairs in batches.
@@
Comprehensive experiments have been conducted, demonstrating that PaDeLLM-NER effectively reduces the number of tokens produced in each sequence, thereby decreasing inference latency. Additionally, it maintains or even enhances prediction quality in both flat and nested NER for English and Chinese languages, compared to existing methods in the literature. To conclude, our contributions are as follows:
@@
  • Comprehensive experiments demonstrate that, in addition to its enhanced prediction speed, PaDeLLM-NER also maintains or surpasses the prediction quality of conventional autoregressive methods, on par with state-of-the-art performance on many NER datasets.
To the best of our knowledge, our techinique stands as a pioneering approach in accelerating NER inference in LLMs by parallel decoding all label-mention pairs. This unique characteristic makes it complementary to other inference acceleration methods such as LLM.int8() Dettmers et al. (2022) and speculative sampling Chen et al. (2023a); Leviathan et al. (2023). Thus, it can be efficiently integrated with these methods.***Code is available at URL masked for anonymous review.
@@ 2.1 Generative Models for NER
Before the era of LLMs, most research approached NER as a sequence labeling task, where each token is assigned a pre-defined tag (e.g., BIO scheme). In this line of work, usually pre-trained transformer-based language models Ma et al. (2020); Liu et al. (2021) is combined with a tailored prediction head to perform a token-level classification, followed by the extraction of identified tokens.
Encouraged by the success of unifying multiple NLP tasks into a single seq2seq paradigm Brown et al. (2020); Lester et al. (2021), especially with the evolution of LLMs Raffel et al. (2020b); Achiam et al. (2023); Touvron et al. (2023a); Yang et al. (2023a), the trend of applying seq2seq models to NER tasks is gaining momentum Xu et al. (2023), with both inputs and outputs being represented as sequences of text Paolini et al. (2020); Lu et al. (2022); Das et al. (2023); Lu et al. (2023a); Wang et al. (2023c). Recently, the focus of work on NER using LLMs has shifted towards zero-shot Xie et al. (2023); Sainz et al. (2023) or few-shot learning Ashok and Lipton (2023); Chen et al. (2023b); Das et al. (2023); Wang et al. (2023b), utilizing in-context learning Chen et al. (2023b); Wang et al. (2023b), self-consistency Wang et al. (2022b); Xie et al. (2023) or learning programming Friedman et al. (2023); Sainz et al. (2023).
Unlike previous studies emphasizing few-shot performance with training-free prompt learning Liu et al. (2023c); Lu et al. (2023b), our work focus on a fully supervised setting. More importantly, our primary objective is to speed up NER inference.
@@ 2.2 Inference Speedup in LLMs
Modern LLMs employ a sequential decoding strategy for token generation, which poses a significant challenge in terms of parallelization, especially as model size and sequence length increase Ning et al. (2023). There is plenty of work in the literature to address this challenge Wang et al. (2021); Frantar et al. (2023); Santilli et al. (2023); Xiao et al. (2023). One line of work falls into training-free category such as introducing extra modules for speculative sampling Chen et al. (2023a); Leviathan et al. (2023). Another approaches explore modifying model architecture to accelerate inference, such as exiting at earlier layer Elbayad et al. (2019); Schuster et al. (2022), or designing entirely different training and inference mechanisms Lan et al. (2023); Yang et al. (2023b); Zhang et al. (2023). Different from previous works, we focus on exploring the inference speedup in LLMs with a focus on the NER task without the change of model architecture or modules.
@@ 3 Method
Figure 1: PaDeLLM-NER training paradigm: texts within frames of the same color represents one training example, where texts inside the solid-line frame are the input, and those inside the dashed-line frame are the output. Italic texts are prompt templates. The “entity type” signifies the label being predicted. The “¡num¿” indicates count of mentions for that label, and “¡mention n¿” refers to the $n^{th}$ mention of a label in the input.
@@ 3.1 Reframing of Instruction Tuning
During reformulation, a single unstructured text containing all label-mention pairs is split into several sequences. Each new sequence’s output includes the count of mentions for a specified label (denoted as “entity type”), followed by the $n^{th}$ mention of that label (denoted as “¡mention n¿”). Note that the count of mentions and their respective indices are represented using corresponding digit tokens from the LLM’s vocabulary. Specifically, if there are no mentions, the model is trained to immediately predict the “¡eos¿” token, bypassing the need to predict mentions.
Therefore, in this example, one original training data is transformed into five new training data entries. These include two for predicting “LOC” (with 2 mentions), one for predicting “MISC” (with 1 mention), one for predicting “PER” (with 1 mention), and one for predicting “ORG” (with 0 mentions, directly predicting “¡eos¿”). Moreover, the number of mentions for each label and the text corresponding to each mention index can be easily obtained from the original ground truth, meaning that the number of new examples depends on the ground truth of that particular example.
@@
where $P\left(t_{i}\mid t_{1},t_{2},\ldots,t_{i-1}\right)$ represents the probability of $i^{th}$ token $t_{i}$ given the sequence of preceding tokens $t_{1},t_{2},\ldots,t_{i-1}$, as predicted by the model. Note that loss calculation begins from the number of mention tokens (i.e., texts enclosed by dashed-line frames). Theoretically, loss from text spans such as “¡mention n¿” could be ignored during this calculation, since they simply prompt the mention’s order, which does not necessarily need to be generated by the model. However, our ablation studies show that ignoring these texts has negligible impact on model performance, a point further discussed in Section 4.3. Therefore, we adhere to the standard instruction tuning procedure.
This reformulation allows the model to focus one label-mention pair at a time, shortening the generated length per sequence. More details are shown in Appendix B.
@@ 3.3 Removal of Duplicate Mentions
To address the issue of duplicate mentions, we suggest employing prediction probability to remove repeated mentions. Specifically, we calculate the prediction probability for each instance of the mention. This is done using the formula: $P=\prod_{i=b}^{e}P(t_{i}|t_{1},t_{2},\ldots,t_{i-1})$ where $b$ represents the starting token index of the mention text, and $e$ denotes the ending token index. Then, for a mention that appears in multiple labels, the mention instance with the highest probability will be preserved. As illustrated in Figure2, “Italy” is categorized as “MISC” with only a 0.61 probability, which is lower than that for “LOC”, resulting in its removal. In practice, the probability of each token can be calculated concurrently with token generation. Consequently, this method enables an efficient and accurate identification of duplicate mentions without incurring additional costs. The effectiveness of this de-duplication approach is further explored in Section 4.3.
@@ 4.1 Setup
We evaluate our method on English and Chinese NER datasets. English datasets include the general domain flat NER CoNLL2003 Tjong Kim Sang and De Meulder (2003), the nested NER ACE2005 Kirkpatrick (2010), and the biomedical nested NER GENIA Ohta et al. (2002). Chinese datasets include four commonly used general domain flat NER benchmarks Resume Zhang and Yang (2018), Weibo Peng and Dredze (2015), MSRA Levow (2006) and Ontonotes 4.0 Pradhan et al. (2013) and two vertical industrial domain flat NER datasets YouKu Jie et al. (2019) and Ecommerce Ding et al. (2019). The statistics of all datasets are shown in Appendix A.
@@
We employ pre-trained version of Llama2-7b Touvron et al. (2023b)***https://huggingface.co/meta-llama/Llama-2-7b and Baichuan2-7b Yang et al. (2023a)***https://huggingface.co/baichuan-inc/Baichuan2-7B-Base as base models for English and Chinese study respectively. Additional implementation details are in Appendix C.
@@
As the primary focus of this work is on reducing inference latency in NER tasks using LLMs, we compare our method, PaDeLLM-NER, with traditional autoregressive approaches. As mentioned in Section 1, the main points of comparison are autoregressive structured output formats used in Paolini et al. (2020); Das et al. (2023) and Lu et al. (2022, 2023a); Wang et al. (2023c), referred to respectively as AutoRegAug and AutoRegStruct, as these are the approaches very close to our system. We reimplemented these methods for both English and Chinese datasets, utilizing the same pre-trained LLMs as in PaDeLLM-NER. More details on the reimplementation are provided in Appendix C. Besides, we compare our approach with other recent state-of-the-art methods, including BINDER Zhang et al. (2022a), Gollie Sainz et al. (2023), and DeepStruct Wang et al. (2022a) for English benchmarks, as well as W2NER Li et al. (2022), NEZHA-BC Zhang et al. (2022b), and SSCNN Zhang and Lu (2023) for Chinese benchmarks, to show PaDeLLM-NER’s efficacy in prediction quality.
@@
Our evaluation encompasses two dimensions: prediction quality and acceleration of NER inference. For assessing prediction quality, in line with Lu et al. (2022); Wang et al. (2023c), we employ the micro F-score.
Regarding inference acceleration, as per Ning et al. (2023), we evaluate using latency (in milliseconds). We record the latency using the following method: start = time.time(); model.generate(); latency = time.time() - start. In PaDeLLM-NER, we add the latency of mention counting and label-mention pair generation as the latency of each sequence. The final latency for the example is determined by the highest latency across sequences, as the user can only obtain the result of an example when the slowest sequence is generated. We conduct experiments three times and use the average result to alleviate the effect of randomness. We also report the average sequence length (tokenized) to clearly demonstrate the extent of sequence length reduction in Appendix D. Evaluations of all models were performed on the same NVIDIA A100 GPU.

### v5

@@ Abstract
In this study, we aim to reduce generation latency for Named Entity Recognition (NER) with Large Language Models (LLMs). The main cause of high latency in LLMs is the sequential decoding process, which autoregressively generates all labels and mentions for NER, significantly increase the sequence length. To this end, we introduce Parallel Decoding in LLM for NER (PaDeLLM-NER), a approach that integrates seamlessly into existing generative model frameworks without necessitating additional modules or architectural modifications. PaDeLLM-NER accelerates decoding by simultaneously generating all mentions at once, i.e., a label-mention pair per sequence. This results in shorter sequences and faster inference. Experiments reveal that PaDeLLM-NER significantly increases inference speed that is 1.76 to 10.22 times faster than the autoregressive approach for both English and Chinese. Concurrently, it maintains the prediction quality as evidenced by the micro F-score that is on par with the state-of-the-art approaches under both zero-shot and supervised setting. All resources are available at https://github.com/GeorgeLuImmortal/PaDeLLM_NER.
footnotetext: Equal contribution. ✉ Corresponding authors.
@@ 1 Introduction
Named Entity Recognition (NER), a fundamental task in Natural Language Processing (NLP), aims to extract structured information from unstructured text data. This includes identifying and categorizing key elements such as Organization, Geopolitical Entity and so on (referred to as “labels”) in inputs, and pairing them with relevant text spans extracted from the text (termed “mentions”). Conventionally, NER tasks are carried out through an extractive paradigm that entails token-level classification and the subsequent extraction of identified tokens [1, 2].
@@
Augmented Language [3, 4] Japan, co-hosts of the World Cup in 2002 and ranked 20th in the world by FIFA, are favourites to regain their title here. [Japan | LOC], co-hosts of the [World Cup | MISC] in 2002 and ranked 20th in the world by [FIFA | ORG], are favourites to regain their title here.
Structured Annotation [5, 6, 7] Cuttitta announced his retirement after the 1995 World Cup , where he took issue with being dropped from the Italy side that faced England in the pool stages. ((PER): (Cuttitta), (MISC): (1995 World Cup), (LOC): (Italy), (LOC): (England), (ORG): (NULL))
@@
Recent advancements in Large Language Models (LLMs) [8, 9, 10, 11, 12, 13] have revolutionized numerous foundational tasks in NLP, including NER tasks [3, 5, 4, 6, 7, 14, 15, 16, 17], through the adoption of a generative paradigm. This paradigm involves instruction-tuning a sequence-to-sequence (seq2seq) model. The model takes a sequence of unstructured text as input and produces a sequence of structured label-mention pairs as output. Generally, the output structured string should be formatted to meet two criteria: (1) it should have a clear and straightforward structure that facilitates post-processing for label and mention extraction, and (2) it needs to be generated fluidly and efficiently from the perspective of language models [18].
@@
In Table 1, we list two typically used autoregressive output formats found in the literature : (1) accommodate original input text to contain label information, which is referred to as “augmented language” [3, 4]; (2) directly using a customized, easily-parsed structured format to output all labels and mentions, which is called “structured annotation” [5, 6, 7]. These formats present certain challenges. For example, augmented language necessitates duplicating all original input text, thereby increasing output length and resulting in inference inefficiency. While structure annotation avoids replicating the entire input, it produces all labels and mentions in an autoregressive manner. This implies that each subsequently generated pair depends on its preceding pairs, and when the number of label-mention pairs is large, it will lead to longer sequences. As demonstrated in Chen et al. [19], Ning et al. [20], high latency in LLMs mainly stems from lengthy sequence generation, we believe that by reducing the length of sequence, a more efficient inference scheme can be provided for NER tasks.
@@
In light of this, we propose Parallel Decoding in LLM for NER (PaDeLLM-NER), a novel approach to accelerate the inference of NER tasks for LLMs. PaDeLLM-NER empowers the model with the capability to predict a single label-mention pair within a single sequence, subsequently aggregating all sequences to generate the final NER outcome. Specifically, in the training phase, we reconstruct the instruction tuning tasks, enabling LLMs to predict the count of mentions for a specific label and to identify the $n^{th}$ mention within the entire input for that label (Figure 1). In the inference phase, LLMs first predict the number of mentions for all labels, then predict all label-mention pairs in parallel (Figure 2). Finally, results from all sequences are aggregated and duplicate mentions across labels are eliminated based on prediction probability. This approach results in a more efficient inference method, producing shorter sequences and enabling parallel decoding label-mention pairs in batches.
@@
Comprehensive experiments have been conducted, demonstrating that PaDeLLM-NER effectively reduces the number of tokens produced in each sequence, thereby decreasing inference latency. Additionally, it maintains or even enhances prediction quality in both flat and nested NER for English and Chinese languages, compared to existing methods in the literature under both zero-shot and supervised setting. To conclude, our contributions are as follows:
@@
  • Comprehensive experiments demonstrate that, in addition to its enhanced prediction speed, PaDeLLM-NER also maintains or surpasses the prediction quality of conventional autoregressive methods, on par with state-of-the-art (SOTA) performance on many NER datasets, including zero-shot as well as the supervised scenarios.
To the best of our knowledge, our technique stands as a pioneering approach in accelerating NER inference in LLMs by parallel decoding all label-mention pairs. This unique characteristic makes it complementary to other inference acceleration methods such as LLM.int8() [21] and speculative sampling [22, 23]. Thus, it can be efficiently integrated with these methods.
@@ 2.1 Generative Models for NER
Before the era of LLMs, most research approached NER as a sequence labeling task, where each token is assigned a pre-defined tag (e.g., BIO scheme). In this line of work, usually pre-trained transformer-based language models [1, 2] is combined with a tailored prediction head to perform a token-level classification, followed by the extraction of identified tokens.
Encouraged by the success of unifying multiple NLP tasks into a single seq2seq paradigm [24, 25], especially with the evolution of LLMs [26, 27, 10, 13], the trend of applying seq2seq models to NER tasks is gaining momentum [28], with both inputs and outputs being represented as sequences of text [3, 5, 4, 6, 7]. Recently, the focus of work on NER using LLMs has shifted towards zero-shot [29, 30] or few-shot learning [31, 32, 4, 18], utilizing in-context learning [32, 18], self-consistency [33, 29] or learning programming [34, 30].
Unlike previous studies emphasizing few-shot performance with training-free prompt learning, our work focus on a fully supervised setting. More importantly, our primary objective is to speed up NER inference.
@@ 2.2 Inference Speedup in LLMs
Modern LLMs employ a sequential decoding strategy for token generation, which poses a significant challenge in terms of parallelization, especially as model size and sequence length increase [20]. There is plenty of work in the literature to address this challenge [35, 36, 37, 38]. One line of work falls into training-free category such as introducing extra modules for speculative sampling [22, 23]. Another approaches explore modifying model architecture to accelerate inference, such as exiting at earlier layer [39, 40], or designing entirely different training and inference mechanisms [41, 42, 43]. Different from previous works, we focus on exploring the inference speedup in LLMs with a focus on the NER task without the change of model architecture or introducing extra modules.
@@ 3 Method
Figure 1: PaDeLLM-NER training paradigm: texts within frames of the same color represents one training example, where texts inside the solid-line frame are the input, and those inside the dashed-line frame are the output. Italic texts are prompt templates. The “entity type” signifies the label being predicted. The “<num>” indicates count of mentions for that label, and “<mention n>” refers to the $n^{th}$ mention of a label in the input.
@@ 3.1 Reframing of Instruction Tuning
During reformulation, a single unstructured text containing all label-mention pairs is split into several sequences. Each new sequence’s output includes the count of mentions for a specified label (denoted as “entity type”), followed by the $n^{th}$ mention of that label (denoted as “<mention n>”). Note that the count of mentions and their respective indices are represented using corresponding digit tokens from the LLM’s vocabulary. Specifically, if there are no mentions, the model is trained to immediately predict the “<eos>” token, bypassing the need to predict mentions.
Therefore, in this example, one original training data is transformed into five new training data entries. These include two for predicting “LOC” (with 2 mentions), one for predicting “MISC” (with 1 mention), one for predicting “PER” (with 1 mention), and one for predicting “ORG” (with 0 mentions, directly predicting “<eos>”). Moreover, the number of mentions for each label and the text corresponding to each mention index can be easily obtained from the original ground truth, meaning that the number of new examples depends on the ground truth of that particular example.
@@
where $P\left(t_{i}\mid t_{1},t_{2},\ldots,t_{i-1}\right)$ represents the probability of $i^{th}$ token $t_{i}$ given the sequence of preceding tokens $t_{1},t_{2},\ldots,t_{i-1}$, as predicted by the model. Note that loss calculation begins from the number of mention tokens (i.e., texts enclosed by dashed-line frames). Theoretically, loss from text spans such as “<mention n>” could be ignored during this calculation, since they simply prompt the mention’s order, which does not necessarily need to be generated by the model. However, our ablation studies show that ignoring these texts has negligible impact on model performance, a point further discussed in Appendix A. Therefore, we adhere to the standard instruction tuning procedure.
This reformulation allows the model to focus one label-mention pair at a time, shortening the generated length per sequence. More details are shown in Appendix C.
@@ 3.3 Removal of Duplicate Mentions
To address the issue of duplicate mentions, we suggest employing prediction probability to remove repeated mentions. Specifically, we calculate the prediction probability for each instance of the mention. This is done using the formula: $P=\prod_{i=b}^{e}P(t_{i}|t_{1},t_{2},\ldots,t_{i-1})$ where $b$ represents the starting token index of the mention text, and $e$ denotes the ending token index. Then, for a mention that appears in multiple labels, the mention instance with the highest probability will be preserved. As illustrated in Figure2, “Italy” is categorized as “MISC” with only a 0.61 probability, which is lower than that for “LOC”, resulting in its removal. In practice, the probability of each token can be calculated concurrently with token generation. Consequently, this method enables an efficient and accurate identification of duplicate mentions without incurring additional costs. The effectiveness of this de-duplication approach is further explored in Appendix A.
@@ 4.1 Setup
The datasets used in our experiments include:
  • Zero-shot Datasets: To align with the methodology proposed by [44], we train PaDeLLM using the Pile-NER dataset [45]. This dataset comprises around 240,000 entities categorized into 13,000 distinct types, derived from the Pile Corpus [46]. The passages in Pile-NER are enhanced through processing with ChatGPT, which facilitates the transparent generation of inherent entities. For assessing the model’s zero-shot capabilities on previously unseen entity categories, following [45, 44, 30] we select two established benchmarks: CrossNER [47] and MIT [48].
  • Supervised Datasets: we evaluate our method on supervised English and Chinese NER datasets. Following [30, 49, 50], English datasets include the general domain flat NER CoNLL2003 [51], the nested NER ACE2005 [52], and the biomedical nested NER GENIA [53]. Following [6, 54, 55], Chinese datasets include four commonly used general domain flat NER benchmarks Resume [56], Weibo [57], MSRA [58] and Ontonotes 4.0 [59] and two vertical industrial domain flat NER datasets YouKu [60] and Ecommerce [61]. The statistics of all datasets are shown in Appendix B.
@@
We employ pre-trained version of Llama2-7b [11]https://huggingface.co/meta-llama/Llama-2-7b and Baichuan2-7b [13]https://huggingface.co/baichuan-inc/Baichuan2-7B-Base as base models for English and Chinese study respectively. Additional implementation details are in Appendix D.
@@
The baseline used in our experiments include:
  • Inference Latency Baselines: As the primary focus of this work is on reducing inference latency in NER tasks using LLMs, we compare our method, PaDeLLM-NER, with traditional autoregressive approaches. As mentioned in Section 1, the main points of comparison are autoregressive structured output formats used in [3, 4] and [5, 6, 7], referred to respectively as AutoRegAug and AutoRegStruct, as these are the approaches very close to our system. We reimplemented all these methods for both English and Chinese datasets, utilizing the same pre-trained LLMs as in PaDeLLM-NER.
  • Zero-shot Baselines: LLMs are known for their generalizability, therefore, following Ding et al. [44], we we also evaluate the zero-shot performance of PaDeLLM. Several most recent SOTA LLM-based approaches are selected as strong baselines as their great generalizability in zero-shot NER scenarios including GoLLIE-7B [30], UniNER-7B [45], GLiNER-L [62], GNER-LLaMA-7B [44].
  • Supervised Baselines: We compare our approach with other recent SOTA supervised approaches, including BINDER [50], Gollie [30], and DeepStruct [49] for English benchmarks, as well as W2NER [63], NEZHA-BC [54], and SSCNN [55] for Chinese benchmarks, to show PaDeLLM-NER’s efficacy in prediction quality.
More details on the re-implementation and model size of each method are provided in Appendix D.
@@
Our evaluation encompasses two dimensions: prediction quality and acceleration of NER inference. For assessing prediction quality, in line with Lu et al. [5], Wang et al. [7], we employ the micro F-score.
Following Ning et al. [20], we evaluate inference speed using latency (in milliseconds). We record the latency with the code: start = time.time(); model.generate(); latency = time.time() - start. In PaDeLLM-NER, we add the latency of mention counting and label-mention pair generation as the latency of each sequence. The final latency for the example is determined by the highest latency across sequences, as the user can only obtain the result of an example when the slowest sequence is generated. We conduct experiments three times and use the average result to alleviate the effect of randomness. We also report the average sequence length (tokenized) to clearly demonstrate the extent of sequence length reduction in Appendix E. Evaluations of all models were performed on the same NVIDIA A100 GPU.